which ordered pair is a solution of the linear system
x+2y=-8 and -4x+y=5

Answer by josgarithmetic(39618)

-4x+y=5 is equivalent to -4x+4x+y=4x+5, y=4x+5.
Substitute into the other given equation:
x+2(4x+5)=-8 and solve for x. What happens? What is y?